OpenStudy (anonymous):

A high school senior applies for admission to two colleges A and B, and suppose that: P (admitted at A) = p1, P (rejected by B) = p2, and P (rejected by at least one, A or B) = p3. i) Calculate the probability that the student is admitted by at least one college. ii) Find the numerical value of the probability in part (i), if p1 = 0.6, p2 = 0.2, and p3= 0.3

OpenStudy (anonymous):

i) here the probility of getting admitted in both college is 1-p3 hence the probability that the student is admitted by at least one college is given by p1+(1-p2) -(1-p3) =p1+p3 -p2

OpenStudy (anonymous):

ii) 0 .6 +0.3 -0.2 =0.7

OpenStudy (anonymous):

Wow thank you!

OpenStudy (anonymous):

welcome dear